# NVS-SQA: Exploring Self-Supervised Quality Representation Learning for Neurally Synthesized Scenes without References

### v1

@@ IV-B Comparison with no-reference quality assessment methods
VIIDEO +$0.222$ ($0.54$) +$0.242$ ($0.45$) +$0.178$ ($0.42$) $-0.412$ ($0.46$) $-0.437$ ($0.35$) $-0.325$ ($0.36$) +$0.143$ ($0.52$) $-0.034$ ($0.56$) +$0.086$ ($0.44$)
Video-BlIINDS +$0.067$ ($0.40$) +$0.107$ ($0.41$) +$0.067$ ($0.34$) +$0.088$ ($0.59$) +$0.206$ ($0.57$) +$0.100$ ($0.50$) +$0.186$ ($0.53$) +$\underline{0.299}$ ($0.50$) +$0.171$ ($0.43$)
FAST-VQA +$0.063$ ($0.56$) +$0.030$ ($0.57$) +$0.022$ ($0.43$) +$0.062$ ($0.65$) $-0.002$ ($0.65$) +$0.075$ ($0.53$) $-0.286$ ($0.45$) $-0.288$ ($0.38$) $-0.257$ ($0.45$)
FasterVQA +$0.104$ ($0.54$) +$0.129$ ($0.62$) +$0.100$ ($0.45$) +$\underline{0.412}$ ($0.46$) +$\underline{0.486}$ ($0.36$) +$\underline{0.375}$ ($0.42$) +$-0.286$ ($0.32$) $-0.160$ ($0.38$) $-0.114$ ($0.30$)
DOVER +$0.189$ ($0.68$) +$0.190$ ($0.66$) +$0.144$ ($0.62$) +$0.112$ ($0.57$) +$0.290$ ($0.56$) +$0.075$ ($0.46$) $-0.129$ ($0.27$) $-0.143$ ($0.37$) $-0.186$ ($0.22$)
DOVER-Mobile +$0.367$ ($0.42$) +$0.449$ ($0.40$) +$0.311$ ($0.39$) +$0.262$ ($0.49$) +$0.268$ ($0.51$) +$0.250$ ($0.40$) $-0.229$ ($0.27$) $-0.175$ ($0.25$) $-0.186$ ($0.24$)
NR-LFQA +$0.168$ ($0.50$) +$0.130$ ($0.53$) +$0.120$ ($0.46$) +$0.052$ ($0.44$) +$0.066$ ($0.42$) +$0.066$ ($0.38$) $-0.102$ ($0.57$) $-0.093$ ($0.52$) $-0.061$ ($0.55$)
Tensor-NLFQ +$0.178$ ($0.48$) +$0.201$ ($0.42$) +$0.159$ ($0.40$) +$0.164$ ($0.35$) +$0.139$ ($0.38$) +$0.071$ ($0.33$) $-0.048$ ($0.58$) $-0.027$ ($0.54$) $-0.081$ ($0.52$)
ALAS-DADS +$0.144$ ($0.46$) +$0.146$ ($0.53$) +$0.156$ ($0.37$) +$0.175$ ($0.54$) +$0.226$ ($0.52$) +$0.125$ ($0.40$) $-0.029$ ($0.19$) +$0.054$ ($0.25$) $-0.029$ ($0.17$)
LFACon +$0.233$ ($0.45$) +$0.182$ ($0.45$) +$0.122$ ($0.42$) +$0.350$ ($0.36$) +$0.385$ ($0.46$) +$0.200$ ($0.35$) +$0.143$ ($0.53$) +$0.212$ ($0.50$) +$\underline{0.186}$ ($0.45$)
Proposed +$\boldsymbol{0.867}$ ($0.08$) +$\boldsymbol{0.812}$ ($0.24$) +$\boldsymbol{0.756}$ ($0.12$) +$\boldsymbol{0.671}$ ($0.12$) +$\boldsymbol{0.647}$ ($0.23$) +$\boldsymbol{0.565}$ ($0.25$) +$\boldsymbol{0.670}$ ($0.09$) +$\boldsymbol{0.669}$ ($0.18$) +$\boldsymbol{0.586}$ ($0.26$)
@@
FSIM +$\underline{0.822}$ ($0.19$) +$\underline{0.785}$ ($0.23$) +$0.711$ ($0.20$) +$0.387$ ($0.49$) +$0.458$ ($0.52$) +$0.350$ ($0.37$) +$0.500$ ($0.41$) +$0.416$ ($0.37$) +$0.429$ ($0.47$)
@@
DSS +$0.789$ ($0.33$) +$0.766$ ($0.27$) +$0.711$ ($0.28$) +$\underline{0.462}$ ($0.48$) +$\underline{0.493}$ ($0.48$) +$\underline{0.425}$ ($0.41$) +$0.257$ ($0.72$) +$0.221$ ($0.74$) +$0.171$ ($0.64$)
HaarPSI +$0.811$ ($0.34$) +$0.747$ ($0.28$) +$\underline{0.756}$ ($0.28$) +$0.362$ ($0.47$) +$0.425$ ($0.52$) +$0.300$ ($0.36$) +$0.343$ ($0.36$) +$0.365$ ($0.32$) +$0.286$ ($0.38$)
@@
LPIPS +$0.722$ ($0.26$) +$0.628$ ($0.27$) +$0.622$ ($0.26$) +$0.187$ ($0.58$) +$0.262$ ($0.62$) +$0.175$ ($0.46$) $-0.300$ ($0.61$) $-0.129$ ($0.65$) $-0.286$ ($0.54$)
DISTS +$0.789$ ($0.17$) +$0.720$ ($0.17$) +$0.667$ ($0.17$) +$0.337$ ($0.50$) +$0.428$ ($0.49$) +$0.300$ ($0.37$) +$0.557$ ($0.28$) +$\underline{0.593}$ ($0.07$) +$0.486$ ($0.30$)
STRRED +$0.800$ ($0.24$) +$0.767$ ($0.30$) +$0.711$ ($0.26$) -- -- -- +$0.543$ ($0.14$) +$0.491$ ($0.31$) +$0.429$ ($0.22$)
VMAF +$0.556$ ($0.50$) +$0.551$ ($0.49$) +$0.511$ ($0.46$) -- -- -- +$0.414$ ($0.13$) +$0.474$ ($0.16$) +$0.286$ ($0.17$)
FovVideoVDP +$0.789$ ($0.33$) +$0.780$ ($0.22$) +$0.733$ ($0.36$) -- -- -- +$\underline{0.657}$ ($0.07$) +$0.472$ ($0.30$) +$\boldsymbol{0.571}$ ($0.14$)
Proposed +$\boldsymbol{0.911}$ ($0.08$) +$\boldsymbol{0.883}$ ($0.11$) +$\boldsymbol{0.822}$ ($0.16$) +$\boldsymbol{0.700}$ ($0.28$) +$\boldsymbol{0.644}$ ($0.39$) +$\boldsymbol{0.625}$ ($0.19$) +$\boldsymbol{0.700}$ ($0.04$) +$\boldsymbol{0.678}$ ($0.10$) +$\boldsymbol{0.571}$ ($0.09$)
@@ IV-C Comparison with full-reference quality assessment methods
Table III delineates the SRCC and PLCC metrics for these full-reference quality assessment methods alongside NVS-SQA, across the Fieldwork, LLFF, and Lab datasets. The table’s concluding row highlights NVS-SQA’s performance improvement over the second-highest scores. Overall, NVS-SQA achieves performance comparable to the evaluated full-reference methods and even demonstrates marginal improvements in some evaluation metrics. Specifically, in the LLFF dataset, NVS-SQA shows an enhancement with a 51.4% increase in the SRCC and an 47.1% rise in the KLCC relative to the top-performing full-reference methods. Moreover, NVS-SQA demonstrates the smallest standard deviation across most of the metrics and datasets, underscoring its stability compared to other methods. These results underscore the exceptional capability of the proposed no-reference method to outperform full-reference methods in the majority of evaluation scenarios, highlighting its potential.
@@ IV-D Ablation study
TABLE IV: Combining all branches outperforms fewer branches, with the auto-weighting method AQB further enhancing performance. This table presents an ablation analysis of the individual and collective efficacy of learning objective branches, comparing manual weighting (MBW) and auto-weighting (AQB). The best results in each column are highlighted in bold to emphasize the best.
@@
AQB $\boldsymbol{0.9111}$ $\boldsymbol{0.8828}$ $\boldsymbol{0.7000}$ $\boldsymbol{0.6441}$ $\boldsymbol{0.7000}$ $\boldsymbol{0.6783}$
[…]
@@
TABLE V: Comparison with the fully-supervised NeRF-NQA. using SRCC and PLCC metrics across the Fieldwork, LLFF, and Lab datasets. Optimal results in each column are highlighted in bold. Given NeRF-NQA was end-to-end trained on perceptual labels, NVS-SQA is also fine-tuned to ensure a fair comparison.
@@
Furthermore, we conducted a comparative analysis between the proposed NVS-SQA framework and our previously introduced fully-supervised method, NeRF-NQA [7]. Given that NeRF-NQA was developed utilizing perceptual labels for training, we accordingly fine-tuned NVS-SQA to ensure an equitable comparison. As delineated in Table V, our method surpasses NeRF-NQA across all evaluation metrics within the three datasets under consideration. Beyond mere performance metrics, NVS-SQA distinguishes itself as an end-to-end learning framework that facilitates straightforward training, in contrast to NeRF-NQA, which does not support end-to-end learning [7]. Specifically, end-to-end learning facilitates the direct transformation of raw data into final outputs, enhancing model efficiency and accuracy by eliminating manual feature engineering [67, 54]. This approach not only simplifies the development process but also improves the scalability and adaptability [67]. Additionally, unlike NeRF-NQA, which relies on external tools such as COLMAP [46]—known for generating noisy sparse points in some cases [68, 69]—NVS-SQA operates independently. This independence further emphasizes its practicality and reliability for the quality assessment of NSS, as it does not depend on external systems that may introduce noise or errors into the process.

### v3

@@ IV-B Comparison with no-reference quality assessment methods
VIIDEO +$0.222$ ($0.54$) +$0.242$ ($0.45$) +$0.178$ ($0.42$) 0.412-0.412 ($0.46$) 0.437-0.437 ($0.35$) 0.325-0.325 ($0.36$) +$0.143$ ($0.52$) 0.034-0.034 ($0.56$) +$0.086$ ($0.44$)
Video-BlIINDS +$0.067$ ($0.40$) +$0.107$ ($0.41$) +$0.067$ ($0.34$) +$0.088$ ($0.59$) +$0.206$ ($0.57$) +$0.100$ ($0.50$) +$0.186$ ($0.53$) $\underline{0.299}$ ($0.50$) +$0.171$ ($0.43$)
FAST-VQA +$0.063$ ($0.56$) +$0.030$ ($0.57$) +$0.022$ ($0.43$) +$0.062$ ($0.65$) 0.002-0.002 ($0.65$) +$0.075$ ($0.53$) 0.286-0.286 ($0.45$) 0.288-0.288 ($0.38$) 0.257-0.257 ($0.45$)
FasterVQA +$0.104$ ($0.54$) +$0.129$ ($0.62$) +$0.100$ ($0.45$) $\underline{0.412}$ ($0.46$) $\underline{0.486}$ ($0.36$) $\underline{0.375}$ ($0.42$) +0.286-0.286 ($0.32$) 0.160-0.160 ($0.38$) 0.114-0.114 ($0.30$)
DOVER +$0.189$ ($0.68$) +$0.190$ ($0.66$) +$0.144$ ($0.62$) +$0.112$ ($0.57$) +$0.290$ ($0.56$) +$0.075$ ($0.46$) 0.129-0.129 ($0.27$) 0.143-0.143 ($0.37$) 0.186-0.186 ($0.22$)
DOVER-Mobile +$0.367$ ($0.42$) +$0.449$ ($0.40$) +$0.311$ ($0.39$) +$0.262$ ($0.49$) +$0.268$ ($0.51$) +$0.250$ ($0.40$) 0.229-0.229 ($0.27$) 0.175-0.175 ($0.25$) 0.186-0.186 ($0.24$)
NR-LFQA +$0.168$ ($0.50$) +$0.130$ ($0.53$) +$0.120$ ($0.46$) +$0.052$ ($0.44$) +$0.066$ ($0.42$) +$0.066$ ($0.38$) 0.102-0.102 ($0.57$) 0.093-0.093 ($0.52$) 0.061-0.061 ($0.55$)
Tensor-NLFQ +$0.178$ ($0.48$) +$0.201$ ($0.42$) +$0.159$ ($0.40$) +$0.164$ ($0.35$) +$0.139$ ($0.38$) +$0.071$ ($0.33$) 0.048-0.048 ($0.58$) 0.027-0.027 ($0.54$) 0.081-0.081 ($0.52$)
ALAS-DADS +$0.144$ ($0.46$) +$0.146$ ($0.53$) +$0.156$ ($0.37$) +$0.175$ ($0.54$) +$0.226$ ($0.52$) +$0.125$ ($0.40$) 0.029-0.029 ($0.19$) +$0.054$ ($0.25$) 0.029-0.029 ($0.17$)
LFACon +$0.233$ ($0.45$) +$0.182$ ($0.45$) +$0.122$ ($0.42$) +$0.350$ ($0.36$) +$0.385$ ($0.46$) +$0.200$ ($0.35$) +$0.143$ ($0.53$) +$0.212$ ($0.50$) $\underline{0.186}$ ($0.45$)
Proposed $\boldsymbol{0.867}$ ($0.08$) $\boldsymbol{0.812}$ ($0.24$) $\boldsymbol{0.756}$ ($0.12$) $\boldsymbol{0.671}$ ($0.12$) $\boldsymbol{0.647}$ ($0.23$) $\boldsymbol{0.565}$ ($0.25$) $\boldsymbol{0.670}$ ($0.09$) $\boldsymbol{0.669}$ ($0.18$) $\boldsymbol{0.586}$ ($0.26$)
@@
FSIM $\underline{0.822}$ ($0.19$) $\underline{0.785}$ ($0.23$) +$0.711$ ($0.20$) +$0.387$ ($0.49$) +$0.458$ ($0.52$) +$0.350$ ($0.37$) +$0.500$ ($0.41$) +$0.416$ ($0.37$) +$0.429$ ($0.47$)
@@
DSS +$0.789$ ($0.33$) +$0.766$ ($0.27$) +$0.711$ ($0.28$) $\underline{0.462}$ ($0.48$) $\underline{0.493}$ ($0.48$) $\underline{0.425}$ ($0.41$) +$0.257$ ($0.72$) +$0.221$ ($0.74$) +$0.171$ ($0.64$)
HaarPSI +$0.811$ ($0.34$) +$0.747$ ($0.28$) $\underline{0.756}$ ($0.28$) +$0.362$ ($0.47$) +$0.425$ ($0.52$) +$0.300$ ($0.36$) +$0.343$ ($0.36$) +$0.365$ ($0.32$) +$0.286$ ($0.38$)
@@
LPIPS +$0.722$ ($0.26$) +$0.628$ ($0.27$) +$0.622$ ($0.26$) +$0.187$ ($0.58$) +$0.262$ ($0.62$) +$0.175$ ($0.46$) 0.300-0.300 ($0.61$) 0.129-0.129 ($0.65$) 0.286-0.286 ($0.54$)
DISTS +$0.789$ ($0.17$) +$0.720$ ($0.17$) +$0.667$ ($0.17$) +$0.337$ ($0.50$) +$0.428$ ($0.49$) +$0.300$ ($0.37$) +$0.557$ ($0.28$) $\underline{0.593}$ ($0.07$) +$0.486$ ($0.30$)
STRRED +$0.800$ ($0.24$) +$0.767$ ($0.30$) +$0.711$ ($0.26$) - - - +$0.543$ ($0.14$) +$0.491$ ($0.31$) +$0.429$ ($0.22$)
VMAF +$0.556$ ($0.50$) +$0.551$ ($0.49$) +$0.511$ ($0.46$) - - - +$0.414$ ($0.13$) +$0.474$ ($0.16$) +$0.286$ ($0.17$)
FovVideoVDP +$0.789$ ($0.33$) +$0.780$ ($0.22$) +$0.733$ ($0.36$) - - - $\underline{0.657}$ ($0.07$) +$0.472$ ($0.30$) $\boldsymbol{0.571}$ ($0.14$)
Proposed $\boldsymbol{0.911}$ ($0.08$) $\boldsymbol{0.883}$ ($0.11$) $\boldsymbol{0.822}$ ($0.16$) $\boldsymbol{0.700}$ ($0.28$) $\boldsymbol{0.644}$ ($0.39$) $\boldsymbol{0.625}$ ($0.19$) $\boldsymbol{0.700}$ ($0.04$) $\boldsymbol{0.678}$ ($0.10$) $\boldsymbol{0.571}$ ($0.09$)
@@ IV-C Comparison with full-reference quality assessment methods
Table III delineates the SRCC and PLCC metrics for these full-reference quality assessment methods alongside NVS-SQA, across the Fieldwork, LLFF, and Lab datasets. The table’s concluding row highlights NVS-SQA’s performance improvement over the second-highest scores. As noted in Table III, dataset LLFF lacks reference video data, making certain full-reference video-based methods such as STRRED, VMAF, and FovVideoVDP inapplicable. Overall, NVS-SQA achieves performance comparable to the evaluated full-reference methods and even demonstrates marginal improvements in some evaluation metrics. Specifically, in the LLFF dataset, NVS-SQA shows an enhancement with a 51.4% increase in the SRCC and an 47.1% rise in the KLCC relative to the top-performing full-reference methods. Moreover, NVS-SQA demonstrates the smallest standard deviation across most of the metrics and datasets, underscoring its stability compared to other methods. These results underscore the exceptional capability of the proposed no-reference method to outperform full-reference methods in the majority of evaluation scenarios, highlighting its potential.
@@ IV-D Ablation study
TABLE IV: Ablation study of individual branch contributions under manual (MBW) and auto-weighting (AQB) schemes. The upper block reports SRCC and PLCC for each single- and pairwise-branch combination using MBW. The lower block shows AQB performance when omitting each branch in turn (w/o IQA, w/o VQA, w/o REP) alongside the fully combined AQB (complete). The best score in each column is highlighted in bold.
@@
AQB w/o IQA $0.7832$ $0.7439$ $0.5827$ $0.5041$ $0.6514$ $0.6422$
AQB w/o VQA $0.8297$ $0.8064$ $0.6028$ $0.5946$ $0.5519$ $0.5233$
AQB w/o REP $0.8463$ $0.8197$ $0.6218$ $0.6135$ $0.6891$ $0.6615$
AQB (complete) $\boldsymbol{0.9111}$ $\boldsymbol{0.8828}$ $\boldsymbol{0.7000}$ $\boldsymbol{0.6441}$ $\boldsymbol{0.7000}$ $\boldsymbol{0.6783}$
We first apply manual weighting MBW (as described in Section 6) with a comprehensive grid search to ascertain the optimal combination of weights for the learning objectives, as specified in Equation 2, within the search space $\{0,\;0.1,\;0.2,\;0.5,\;1,\;1.5,\;2\}$. The search culminated in the identification of the most effective weight combination: $\{\lambda_{IQA}:1.5,\;\lambda_{VQA}:1,\;\lambda_{REP}:0.2\}$. We then applied the auto-weighting method AQB (introduced in Section 6) to evaluate its effectiveness. Table IV presents the results of ablation studies, assessing the individual and collective contributions of each learning branch, as well as the impact of AQB. The studies conducted on Fieldwork, LLFF, and Lab demonstrate that combining all branches in MBW outperforms using fewer branches, with AQB further enhancing performance. Moreover, we conduct an extended ablation study to analyze the contribution of each guidance branch (IQA, VQA, and REP) under both manual weighting (MBW) and the proposed adaptive scheme (AQB). As shown in Table IV, we report results for each individual branch, their combinations, and the complete model. In addition, we include AQB-based ablations that selectively remove one branch at a time. These results show that the IQA branch is critical, with a marked performance degradation observed upon its removal. The VQA branch also plays a significant role, particularly in capturing angular and temporal coherence. Meanwhile, the REP branch provides complementary quality features not captured by IQA or VQA, further enhancing overall performance. This comprehensive analysis further validates the importance of each component in our multi-branch design and the effectiveness of AQB in leveraging them optimally. Fig. 8 qualitatively illustrates the influence of each guidance branch on the model’s prediction preference. When any of the three branches is removed, NVS-SQA tends to assign higher quality scores to perceptually inferior scenes, such as those exhibiting blurring, motion instability, or structural misalignment, revealing a reduced sensitivity to the corresponding perceptual dimension.
Scene [Room] Synthesized by NVS methods (NeX) and (GNT-C)
Predicted Preference w/o IQA Branch Predicted Preference (with IQA Branch)
Refer to caption Refer to caption
Scene [Whale] Synthesized by NVS methods (GNT-C) and (IBRNet-S)
Predicted Preference w/o VQA Branch Predicted Preference (with VQA Branch)
Refer to caption Refer to caption
Scene [Leaves] Synthesized by NVS methods (NeRF) and (NeX)
Predicted Preference w/o REP Branch Predicted Preference (with REP Branch)
Refer to caption Refer to caption
Figure 8: Qualitative illustration of the effects of individual guidance branches (IQA, VQA, REP) on the prediction preference of the proposed NVS-SQA model. Each row shows two synthesized scenes generated by different NVS methods. The left and right columns indicate which scene is preferred (i.e., predicted as higher quality) by NVS-SQA when trained without or with a specific guidance branch. For example, when the IQA branch is omitted, the model tends to favor scenes with blurred or texture-inconsistent regions, while inclusion of the IQA branch shifts preference toward perceptually sharper results. Similar trends are observed for the VQA and REP branches, where their inclusion improves sensitivity to temporal consistency and structural integrity, respectively. These results demonstrate that each guidance branch contributes a distinct and complementary perceptual cue to the overall quality representation (zooming in for a clearer view).
To evaluate the sensitivity of NVS-SQA to training hyperparameters, we varied the number of training epochs and batch sizes while keeping all other factors constant, including the architecture, AQB loss function, and unlabeled training data. The results, presented in Table V, indicate that NVS-SQA maintains robust performance across a range of batch sizes when the number of training epochs is sufficient. This suggests that the model is not overly sensitive to moderate changes in these hyperparameters.
TABLE V: Sensitivity analysis of training hyperparameters. This table reports the SRCC, PLCC, and KRCC values under different combinations of epochs and batch sizes using the Fieldwork dataset.
Epochs Batch Size SRCC $\uparrow$ PLCC $\uparrow$ KRCC $\uparrow$
100 16 $0.8352$ $0.7823$ $0.7538$
200 16 0.9110 0.8831 0.8226
300 16 $0.9034$ $0.8707$ $0.8142$
200 8 $0.8796$ $0.8549$ $0.7931$
200 32 $0.9012$ $0.8684$ $0.8107$
To evaluate the contribution of our proposed contrastive pair preparation, we introduced a baseline that removes the entire pairing mechanism. Specifically, this variant samples NSS pairs randomly across the dataset without ensuring semantic consistency, view replacement, or controlled distortion. The results can be found in the first and last row of Table VI. The absence of these constraints substantially degrades the model’s ability to learn meaningful quality representations: for example, SRCC on the Fieldwork dataset drops from 0.911 to 0.745. This confirms that our contrastive pair preparation strategy provides essential structure for learning discriminative quality embeddings.
In addition, we explored several alternatives to our Transformer-based angular fusion design within AdaptiSceneNet. We replaced the Transformer module with convolution-only layers, GRUs, and LSTMs, respectively. The results can be found in Table VI. While both GRU and LSTM architectures offered improvements over CNN-only designs, they consistently underperformed compared to our Transformer-based approach. For example, on the Fieldwork dataset, SRCC improved from 0.714 (CNN-only) to 0.820 (CNN+LSTM), but remained below the 0.911 achieved by our full model. These results validate the advantage of using self-attention to model long-range dependencies across views, which is critical for capturing angular consistency in NSS.
TABLE VI: Ablation on contrastive pair preparation and angular fusion design. This table reports SRCC values on the Fieldwork, LLFF, and Lab datasets using different architectural and training configurations.
Model Variant Fieldwork $\uparrow$ LLFF $\uparrow$ Lab $\uparrow$
w/o Contrastive Pair Preparation $0.7452$ $0.6651$ $0.6874$
CNN-only $0.7147$ $0.6303$ $0.6182$
CNN + GRU $0.7856$ $0.7189$ $0.7130$
CNN + LSTM $0.8203$ $0.7458$ $0.7507$
CNN + Transformer (Ours) 0.9110 0.8831 0.8226
@@
TABLE VII: Comparison with the fully-supervised NeRF-NQA. using SRCC and PLCC metrics across the Fieldwork, LLFF, and Lab datasets. Optimal results in each column are highlighted in bold. Given NeRF-NQA was end-to-end trained on perceptual labels, NVS-SQA is also fine-tuned to ensure a fair comparison.
@@
Furthermore, we conducted a comparative analysis between the proposed NVS-SQA framework and our previously introduced fully-supervised method, NeRF-NQA [7]. Given that NeRF-NQA was developed utilizing perceptual labels for training, we accordingly fine-tuned NVS-SQA to ensure an equitable comparison. As delineated in Table VII, our method surpasses NeRF-NQA across all evaluation metrics within the three datasets under consideration. Beyond mere performance metrics, NVS-SQA distinguishes itself as an end-to-end learning framework that facilitates straightforward training, in contrast to NeRF-NQA, which does not support end-to-end learning [7]. Specifically, end-to-end learning facilitates the direct transformation of raw data into final outputs, enhancing model efficiency and accuracy by eliminating manual feature engineering [67, 54]. This approach not only simplifies the development process but also improves the scalability and adaptability [67]. Additionally, unlike NeRF-NQA, which relies on external tools such as COLMAP [46]—known for generating noisy sparse points in some cases [68, 69]—NVS-SQA operates independently. This independence further emphasizes its practicality and reliability for the quality assessment of NSS, as it does not depend on external systems that may introduce noise or errors into the process.